# Applying statistical learning theory to deep learning

### v1

@@ Abstract
Although statistical learning theory provides a robust framework to understand supervised learning, many theoretical aspects of deep learning remain unclear, in particular how different architectures may lead to inductive bias when trained using gradient based methods. The goal of these lectures is to provide an overview of some of the main questions that arise when attempting to understand deep learning from a learning theory perspective. After a brief reminder on statistical learning theory and stochastic optimization, we discuss implicit bias in the context of benign overfitting. We then move to a general description of the mirror descent algorithm, showing how we may go back and forth between a parameter space and the corresponding function space for a given learning problem, as well as how the geometry of the learning problem may be represented by a metric tensor. Building on this framework, we provide a detailed study of the implicit bias of gradient descent on linear diagonal networks for various regression tasks, showing how the loss function, scale of parameters at initialization and depth of the network may lead to various forms of implicit bias, in particular transitioning between kernel or feature learning.
@@ 1.1 Preamble
We start by giving you Nati’s view of supervised learning and statistical learning theory, and the main forces in learning. We’ll mostly focus on the concept of inductive bias. We’ll be starting by trying to explain what is inductive bias, or rather, what I mean by inductive bias, as this term is hard to define mathematically. We’ll incorporate this idea of inductive bias in the introduction of statistical learning theory, and try to understand how it applies to deep learning. We’ll explore how deep learning fits into the approach we developed to understand learning theory in the past five decades. The main goal for today is to mention what we need to understand (what questions we need to answer), and what notions we need to rethink.
@@ 1.2 Inductive bias in supervised learning
The goal of supervised learning is to find a predictor, a mapping $h$ from inputs or instances $\mathcal{X}$ (e.g. images, sentences) to labels $\mathcal{Y}$ (e.g. classes). In the simplest setting we’ll just think of $y\in\{\pm 1\}$. The goal is to find a predictor that has a small generalization loss $L(h)$. Crucially, we measure the loss with respect to a source distribution $\mathcal{D}$ and our goal is to find $h$ which in expectation over this distribution has small loss, meaning its predictions tend to be correct.
Supervised learning: find $\mbox{find }h:\mathcal{X}\rightarrow\mathcal{Y}\mbox{ with small {generalization error}}$
@@
Machine learning designs this predictor $h$, not based on knowledge of the population $\mathcal{D}$ but rather based on an IID sample from that population. We try to find good learning rules that produce a predictor with small error for any population.
@@
$A:\ S\rightarrow h\quad\mbox{(i.e. $A:(\mathcal{X}\times\mathcal{Y})^{*}\rightarrow\mathcal{Y}^{\mathcal{X}}$)}$
Unfortunately, this is impossible. The ‘no free lunch’ theorem tells us small generalization error requires knowledge about the population. For any learning rule, there exists some distribution $\mathcal{D}$ (that is, some reality) for which the learning rule yields an expected error that is tantamount to randomly guessing the answer (e.g. $1/2$ for binary classes). More formally, for any $A,m$ there exists $\mathcal{D}$ such that $\exists_{h^{*}}L(h^{*})=0$ but
$\mathbb{E}_{S\sim\mathcal{D}^{m}}[L(A(S))]\geq\frac{1}{2}-\frac{m}{2|\mathcal{X}|}$
This is true not only for independent $x,y$, but also if there exists a deterministic relation $y(x)$, so that a predictor does exist. The supposed improvement over $1/2$ is proportional to the size of the dataset $m/|\mathcal{X}|$ is due to memorization of that dataset, and vanishes when the population size is large.
Thus, learning is impossible without assuming anything. This is where inductive bias becomes an essential part of learning. We assume that some realities (populations $\mathcal{D}$) are unlikely, and design the learning rule $A$ to work for the more likely realities, e.g. by preferring certain models $h(x)$ over others. More practically, we assume reality $\mathcal{D}$ has a certain property which ensures the learning rule $A$ has good generalization error. Typically, we assume that there exist models $h(x)$ in some class $\mathcal{H}$, or with low “complexity” $c(h)$ such that it has low generalization error $L(h)$. An example are models where the output $y$ changes smoothly with the input $x$. Another example is ridge regression, that prefers linear models (with small norms of the weights).
A flat inductive bias embodies the assumption that some realities are possible and others are not, $\exists{h^{*}}\in\mathcal{H}$ with low $L(h^{*})$. If we make this assumption, we know what is the best learning rule for supervised learning, which is empirical risk minimization:
$ERM_{\mathcal{H}}(S)=\hat{h}=\arg\min_{h\in\mathcal{H}}L_{S}(h)$
with $L_{S}$ the empirical, or training loss over our sample $S$ of size $m$:
@@
For this learning rule, we can guarantee an upper bound on the generalization error. If the best model in our class is $h^{*}$, the error of the predictor achieved by the ERM learning rule is
@@
We see that the error is larger than the best error in our class $\mathcal{H}$ by a term that scales with a measure of its capacity (colloquially, the ‘amount’ of models in the class), divided by the number of samples. In other words, the number of training examples required to learn $h^{*}$ scales with the capacity of class $\mathcal{H}$.
What is this capacity? For binary classification, the capacity is the Vapnik-Chrvonenkis (VC) dimension of the class, $capacity(\mathcal{H})=VCdim(\mathcal{H})$. The VC dimension is the largest number of points $D$ that can be labeled (by models $h\in\mathcal{H}$) in every possible way. This is quite natural. A model class with high VC dimension (i.e. that contains predictors allowing any possible labelling of the set), does not have any inductive bias, so that learning is impossible (no free lunch). Learning becomes possible when the model class can be falsified, and the number of samples needed for learning is the number required to falsify this assumption on the model class $\mathcal{H}$. For linear classifiers over $d$ features, $VCdim(\mathcal{H})=d$. In fact, if the model class $\mathcal{H}$ can be parameterized with $d$ parameters, the VC dimension is usually $VCdim(\mathcal{H})=\tilde{O}(d)$. It is always true that the VC dimension is bounded by the logarithm of the cardinality of $\mathcal{H}$:
@@
Thus we expect that if we encode the parameters of our model with a fixed number of bits, the VC dimension of the model scales with the number of parameters. Another way to produce model classes with finite capacity is employ regularizers in our learning rule. For example, it can be shown that for linear predictors with norm $||w||_{2}\leq B$ (with logistic loss and normalized data), $capacity(\mathcal{H})=B^{2}$.
Looking back at (*)(*)( * ), we see that learning, and machine learning in particular, requires model (hypothesis) classes $\mathcal{H}$ that are expressive enough to approximate reality well (contain $h^{*}$ with low generalization error), but also have a small enough capacity to allow for good generalization. The approximation error is defined by the error of the best model in our class $h^{*}$, and the estimation error is the excess over it, as the learning rule can choose a different, worse model given the empirical data.
Usually however, our learning protocols do not represent a flat inductive bias over some model class. We often think in terms of a complexity measure $c:\ \mathcal{Y}^{\mathcal{X}}\rightarrow[0,\infty]$, which is any ordering of predictors $h$. Some measures of complexity include the degree of polynomials is our model class, assumptions of sparsity or low norm $||w||$. The associated inductive bias is that $\exists h^{*}$ with low complexity $c(h^{*})$ and low error $L(h^{*})$. This inductive bias suggests another learning rule, structural risk minimization:
$SRM_{\mathcal{H}}(S)=\arg\min_{h\in\mathcal{H}}L_{S}(h),c(h)$
This learning rule attempts to minimize two functions, which naturally introduces a trade-off between them. At best, the learning rule achieves a predictor that sits on the Pareto frontier that trades-off generalization error and complexity. Any predictor on that line cannot improve either the error or complexity without worsening the other. You can get to this frontier by considering a regularization path, i.e. minimizing $L_{S}(h)+\lambda c(h)$, and varying the regularization amplitude $\lambda$ in the range $[0,\infty]$. Equivalently, one can attempt to minimize the error $L_{S}$ such that $c(h)\leq B$. Note that this learning rule retrieves a multitude of candidates along the Pareto frontier. We can choose the best of them according to their performance on a cross validation sample.
For the SRM learning rule, we get a similar guarantee to (*)(*)( * ) on the generalization error, although
@@ 1.3 Inductive bias in deep learning
Deep learning is learning with a particular inductive bias, a flat hypothesis class in the form of a feed-forward neural network. A feed-forward neural network (Fig. 1) is described by a directed graph $G(V,E)$ with nodes (neurons) indexed by vertices $V$. These nodes are subdivided into three types:
@@
The network also has an activation function $\sigma:\mathbb{N}\rightarrow\mathbb{N}$ that describes the non-linearity of the neural network; a popular choice is the rectified linear unit (ReLU), $\sigma_{ReLU}=[z]_{+}$. Finally, the edges of the network, each connecting $2$ nodes $u,v$, are called weights $w[u\rightarrow v]$ for each edge $u\rightarrow v\in E$. A choice of architecture, weights and activation function uniquely describe a predictor function $h_{w}(x)$. These models were historically developed by McCulloch and Pitts to describe computation in the 1940’s [19]. They where able to show these models can perform complex computations.
@@
As noted before, the capacity is roughly given by the number of learned parameters, which here is the number of edges $|E|$. The VC dimension of these networks with threshold activation is $VCdim(\mathcal{H}_{G(V,E),sign})=O\left(|E|\log|E|\right)$. However, for other activation functions, it is actually possible to get a capacity which is much higher than the number of parameters. For example, the VC dimension of these networks with sine activation is infinite, even for a single hidden node. We do not use these kind of activation functions. More useful activation functions are, for example the sigmoid function $\sigma(z)=(1+\exp(z))^{-1}$, whose VC dimensions is bounded by $VCdim(\mathcal{H}_{G(V,E),sigmoid})\leq O\left(|E|^{4}\right)$, or ReLU function, for which $VCdim(\mathcal{H}_{G(V,E),ReLU})\leq O\left(|E|\log|E|l\right)$, with $l$ the network depth. One can limit the capacity by discretizing the weights, e.g. if $w\in[-B,...,B]$, $VCdim\leq 2|E|B$. As we’ve seen, the fact that these network models can have a large capacity is not necessarily good, because capacity comes at the expense of inductive bias.
What about expressivity? Feed-forward neural networks can represent any logical gate, and thus any function over $\mathcal{X}=\{\pm 1\}^{d}$ (as proved by Turing). Define the class $CIRCUIT_{n}[depth,size]$ as all functions $f:\{\pm 1\}^{n}\rightarrow\{0,1\}$ that can be implemented with at most $size$ AND, OR and NOT gates, and longest path from input to output at most $depth$. We know that circuits can represent any function, but only if we are allowed to select an appropriate gate architecture. In neural networks, we keep the architecture fixed and only vary the weights.
@@
This can be done easily, if we choose the weights of each edge to be ±$\pm 1$ if the edge is connected in the circuit (with /o without a NOT gate in between), 00 otherwise. The bias terms are chosen $fan\_in-1$ for AND gates, $1-fan\_in$ for OR gates. The weights essentially describe which wires exist in the circuit. Thus neural networks can represent any binary function.
More generally, we have a universal representation theorem: Any continuous function $f:[0,1]^{d}\rightarrow\mathbb{N}$ can be approximated to within $\epsilon$ by a feed-forward network with sigmoidal (or almost any other) activation function and a single hidden layer. This shows that as a model class, feed-forward neural nets are extremely expressive and can represent any reality. However, representing functions may require huge networks, e.g. with layer widths exponential in $d$. The relevant question is not what a network of arbitrary size can represent, but what small networks can represent.
@@
Interestingly, a feed-forward neural network can represent any time $T$ computable function with network of size $\tilde{O}(T)$ using a depth-$T$ network. This is true since anything computable in time $T$ is also computable by a logical circuit of size $\tilde{O}(T)$. This realization has broad implications for machine learning.
Machine learning is an engineering paradigm (of being lazy): Use data and examples, instead of expert knowledge and tedious programming, to automatically create efficient systems that solve complex tasks. Therefore, we only care about a model (predictor) $h$ if it can be implemented efficiently. A good learned model only needs to compete with a programmer, producing results that are at least as good as a programmed model in a competitive (model evaluation) time.
In this case we have a free lunch: the model class $TIME_{T}$ all functions computable by at most time $T$, has capacity $O\left(T\right)$, and hence learnable with $O\left(T\right)$ parameters (e.g. using ERM). Even better: the model class $PROG_{T}$, all functions programmable up to length of code $T$, also has capacity $O\left(T\right)$. This is relatively clear, because the length $T$ bounds the number of bits needed to represent all these functions. Unfortunately, ERM with respect to $PROG_{T}$ is uncomputable. Modified ERM for $TIME_{T}$ (truncating exec. time) can be computed, but is NP-complete. If $P=NP$, we can have universal learning (free lunch). If $P\neq NP$, i.e. if there exist one-way functions that are easy to compute but hard to learn, there is no poly-time algorithm for ERM over $TIME_{T}$.
@@
Going back to neural networks, we completely understand them from a statistical perspective (in terms of capacity and expressivity). The problem with them relates to computation; computing the ERM for feed-forward neural nets is a non-convex optimization problem, and no known algorithm is guaranteed to work. We know that learning in neural nets, even in the simplest cases ($2$-hidden units in one hidden layer), is NP-hard. Even if reality is well-approximated by a small neural net, and you tried optimizing a larger neural net (which has more degrees of freedom), optimization is easy but ERM is still NP-hard. Unfortunately, there is nothing one can do to efficiently solve this computational problem, which is essential to neural nets precisely because of their expressive power (computational no-free lunch). Even if a function is exactly representable with single hidden layer and $\Theta(\log d)$ nodes, even with no noise, and even if we take a much larger network or use any other method when learning: no poly-time algorithm can ensure better-than-chance prediction, see e.g. [14, 10].
And nevertheless, deep learning does work! We have seen that from a statistical and computational perspectives, performing ERM on short programs (or short runtime programs) and learning with deep networks is equivalent. Both approaches are universal and approximate reality with reasonable sample complexity. They are both NP-hard, and provably hard to learn with any learning rule (subject to cryptographic assumptions). However there is no practical way to optimize over short programs, as e.g. there is no practical local search over programs. In contrast, deep neural nets are often easy to optimize; they are continuous models, amenable to local search (gradient descent, SGD), and enjoy much empirical success. In the worst case, deep learning is provably impossible, and yet, we are constantly reminded that deep learning is possible. There is a certain magical property of reality that makes feed-forward neural networks, and we have just started to scratch the surface of what that property is. However, we know for sure what it isn’t: it is not the property that reality is well approximated by neural networks.

### v2

@@ Abstract
Although statistical learning theory provides a robust framework to understand supervised learning, many theoretical aspects of deep learning remain unclear, in particular how different architectures may lead to inductive bias when trained using gradient based methods. The goal of these lectures is to provide an overview of some of the main questions that arise when attempting to understand deep learning from a learning theory perspective. After a brief reminder on statistical learning theory and stochastic optimization, we discuss implicit bias in the context of benign overfitting. We then move to a general description of the mirror descent algorithm, showing how we may go back and forth between a parameter space and the corresponding function space for a given learning problem, as well as how the geometry of the learning problem may be represented by a metric tensor. Building on this framework, we provide a detailed study of the implicit bias of gradient descent on linear diagonal networks for various regression tasks, showing how the loss function, scale of parameters at initialization and depth of the network may lead to various forms of implicit bias, in particular transitioning between kernel and feature learning regimes.
@@ 1.1 Preamble
The purpose of this first lecture is to provide a mathematical framework that allows to introduce the notion of inductive (or implicit) bias in a clear way. We will see that this notion is already present in the usual formulation of supervised learning, as it implicitly appears when preferring certain classes of model over others in empirical risk minimization procedures. While the statistical aspect of supervised learning can be understood by measuring the complexity of classes of predictors and introducing an associated notion of inductive bias, the computational aspect also requires an assumption of implementability. This is especially relevant in the context of deep learning, as the expressivity of neural networks can seemingly lead to the representation and learning of arbitrary functions. This will naturally lead us to the notion of computationally efficient learning rule, and to turn our attention to the implicit bias associated with the optimization aspect of supervised learning.
@@ 1.2 Inductive bias in supervised learning
The main idea of supervised learning is to find a predictor, i.e. a mapping $h$ belonging to some function space $\mathcal{H}$, from inputs or instances $\mathcal{X}$ (e.g. images, sentences) to labels $\mathcal{Y}$ (e.g. classes) in order to predict the labels of new instances. In the simplest setting we’ll just think of $y\in\{\pm 1\}$. To do so, we introduce a loss function $loss:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{+}$ which quantifies the error made by our predictor for a given labeled example $(x,y)\in\mathcal{X}\times\mathcal{Y}$. The goal is to find a predictor that has a small generalization loss $L(h)$ defined as the expected value of our loss function $loss$ with respect to a source distribution $\mathcal{D}$. Intuitively, this amounts to find $h$ such that our predictions tend to reproduce the hidden joint distribution. This is formalized in the following definition :
Supervised learning: find $\mbox{find }h:\mathcal{X}\rightarrow\mathcal{Y}\mbox{ with small {generalization error}, defined as},$
@@
A central aspect of machine learning is to design this predictor $h$, not based on knowledge of the population $\mathcal{D}$, but rather based on an IID sample $\mathcal{S}$ (to which we have access to through an experiment, a dataset, …) from that population. We attempt to find a good learning rule, i.e. a mapping $A:\mathcal{S}\to\mathcal{H}$, that produces a predictor with small error for any population.
@@
$A:\ S\rightarrow h\quad\mbox{(i.e. $A:(\mathcal{X}\times\mathcal{Y})\rightarrow\mathcal{Y}^{\mathcal{X}}$)}$
Unfortunately, this is impossible. The ‘no free lunch’ theorem , a terminology originally introduced in [52], tells us that small generalization error requires knowledge about the population. For any learning rule, there exists some distribution $\mathcal{D}$ (that is, some reality) for which the learning rule yields an expected error that is tantamount to randomly guessing the answer (e.g. $1/2$ for binary classification). More formally, for any learning rule $A$ based on an IID sample $\mathcal{S}$ of size $m$, there exists a distribution $\mathcal{D}$ such that there exists a predictor $h^{*}$ verifying $L(h^{*})=0$, but
$\mathbb{E}_{S\sim\mathcal{D}^{m}}[L(A(S))]\geq\frac{1}{2}-\frac{m}{2|\mathcal{X}|},$
where $|\mathcal{X}|$ designates the cardinality of the input space $\mathcal{X}$, which may be infinite. This is true not only for independent $x,y$, but also if there exists a deterministic relation $y(x)$, so that a predictor does exist. The supposed improvement over $1/2$ is proportional to the size of the dataset $m/|\mathcal{X}|$ and is due to memorization of that dataset, and vanishes when the population size is large.
Thus, learning is impossible without assuming anything. This is where inductive bias becomes an essential part of learning. We assume that some realities (populations $\mathcal{D}$) are unlikely, and design the learning rule $A$ to work for the more likely realities, e.g. by preferring certain models $h(x)$ over others. More practically, we assume that the reality $\mathcal{D}$ has a certain property which ensures the learning rule $A$ will have a good generalization error. Typically, we assume that there exist models $h(x)$ in some class $\mathcal{H}$, or with low “complexity”, denoted $c(h)$, such that it has low generalization error $L(h)$. An example are models where the output $y$ changes smoothly with the input $x$, where the complexity of the model can be captured by its total variation or an appropriate Sobolev norm. Another example is ridge regression, that prefers linear models, in which case the complexity will be measured with the norm of the weights of the predictor.
A flat inductive bias embodies the assumption that some realities are possible and others are not : $\exists{h^{*}}\in\mathcal{H}$ with low $L(h^{*})$. If we make this assumption, we know what is the best learning rule for supervised learning, which is empirical risk minimization:
$ERM_{\mathcal{H}}(S)=\hat{h}=\arg\min_{h\in\mathcal{H}}L_{S}(h),$
with $L_{S}$ the empirical loss, or training loss over our sample $S$ of size $m$:
@@
For this learning rule, we can guarantee an upper bound on the generalization error based on the capacity of the hypothesis class $\mathcal{H}$, which quantifies the complexity of our model. If the best model in our class is $h^{*}$, the error of the predictor achieved by the ERM learning rule is
@@
where the quantity $\mathcal{R}_{m}(\mathcal{H})$ represents the Rademacher complexity of the class $\mathcal{H}$ and will be studied more carefully in Lecture 3. For now, let us simply note that the error is larger than the best error in our class $\mathcal{H}$ by a term that scales with a measure of its capacity (colloquially, the ‘amount’ of models in the class), divided by the number of samples. In other words, the number of training examples required to learn $h^{*}$ scales with the capacity of the class $\mathcal{H}$.
Let us study this notion of capacity more carefully. For binary classification, the capacity is the Vapnik-Chervonenkis (VC) [50, 49] dimension of the class, $capacity(\mathcal{H})=VCdim(\mathcal{H})$. The VC dimension is the largest number of points $D$ that can be labeled by models $h\in\mathcal{H}$ in every possible way. Thus it quantifies the ability of the models in our hypothesis class $\mathcal{H}$ to fit an arbitrarily labeled dataset. This is quite natural. A model class with high VC dimension (i.e. that contains predictors allowing any possible labelling of the set), does not have any inductive bias, so that learning is impossible (no free lunch). Learning becomes possible when the model class can be falsified, and the number of samples needed for learning is the number required to falsify this assumption on the model class $\mathcal{H}$. For linear classifiers over $d$ features, $VCdim(\mathcal{H})=d$. In fact, if the model class $\mathcal{H}$ can be parameterized with $d$ parameters, the VC dimension is usually $VCdim(\mathcal{H})=\tilde{O}(d)$. It is always true that the VC dimension is bounded by the logarithm of the cardinality of $\mathcal{H}$:
@@
Thus we expect that if we encode the parameters of our model with a fixed number of bits, the VC dimension of the model scales with the number of parameters. Another way to produce model classes with finite capacity is to employ regularizers in our learning rule, explicitly penalizing models with high complexity. For example, it can be shown that for linear predictors with norm $||w||_{2}\leq B$ (with logistic loss and normalized data), $capacity(\mathcal{H})=B^{2}$. More detail on VC dimension and related notions can be found in e.g. [49, 13].
Looking back at ()(*)( ∗ ), we see that learning, and machine learning in particular, requires model (hypothesis) classes $\mathcal{H}$ that are expressive enough to approximate reality well (contain $h^{*}$ with low generalization error), but also have a small enough capacity to allow for good generalization. The approximation error is defined by the error of the best model in our class $h^{*}$, and the estimation error is the excess over it, as the learning rule can choose a different, worse model given the empirical data.
Usually however, our learning protocols do not represent a flat inductive bias over some model class. We often think in terms of a complexity measure $c:\ \mathcal{Y}^{\mathcal{X}}\rightarrow[0,\infty]$, which is formally any ordering of predictors $h$. Some measures of complexity include for instance the degree of polynomials is our model class, the cardinality of active weights of the predictor (i.e. an assumption of sparsity) or a low given norm of the weights of our predictor $||w||$. The associated inductive bias is that $\exists h^{*}$ with low complexity $c(h^{*})$ and low error $L(h^{*})$. This inductive bias suggests another learning rule, structural risk minimization:
$SRM_{\mathcal{H}}(S)=\arg\min_{h\in\mathcal{H}}L_{S}(h),c(h).$
This learning rule attempts to minimize two functions, which naturally introduces a trade-off between them. At best, the learning rule achieves a predictor that sits on the Pareto frontier that trades off generalization error and complexity. Any predictor on that line cannot improve either the error or complexity without worsening the other. It is possible to achieve this frontier by considering a regularization path, i.e. minimizing $L_{S}(h)+\lambda c(h)$, and varying the regularization amplitude $\lambda$ in the range $[0,\infty]$. Equivalently, one can attempt to minimize the error $L_{S}$ such that $c(h)\leq B$ (in which case the aforementioned $\lambda$ can be understood as a Lagrange multiplier). Note that this learning rule retrieves a multitude of candidates along the Pareto frontier. We can choose the best of them, for instance, according to their performance on a cross validation sample.
For the SRM learning rule, we get a similar guarantee to ()(*)( ∗ ) on the generalization error, although our hypothesis class $\mathcal{H}$ is now reduced to functions with low complexity, and in particular for the optimal predictor $h^{*}$, we have the bound
@@ 1.3 Inductive bias in deep learning
Deep learning is learning with a particular inductive bias, a flat hypothesis class in the form of a feed-forward neural network. For more information on the history of neural networks and their applications, the interested reader may consult the textbook [28]. A feed-forward neural network (Fig. 1) is described by a directed graph $G(V,E)$ with nodes (neurons) indexed by vertices $V$. These nodes are subdivided into three types:
@@
The network also has an activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ that describes the non-linearity of the neural network; a popular choice is the rectified linear unit (ReLU), $\sigma_{ReLU}(z)=[z]_{+}=\max(z,0)$. Finally, the edges of the network, each connecting $2$ nodes $u,v$, are called weights $w[u\rightarrow v]$ for each edge $u\rightarrow v\in E$. A choice of architecture, weights and activation function uniquely describe a predictor function $h_{w}(x)$. These models were historically developed by McCulloch and Pitts to describe logical calculus related to nervous activity [33]. They were able to show that these models can perform complex computations, with the complexity measure directly related to the cardinality of the net.
@@
As noted before, the capacity is roughly given by the number of learned parameters, which here is the number of edges $|E|$. The VC dimension of these networks with threshold activation (ReLu) is $VCdim(\mathcal{H}_{G(V,E),sign})=O\left(|E|\log|E|\right)$. However, for other activation functions, it is actually possible to get a capacity which is much higher than the number of parameters. For example, the VC dimension of these networks with sine activation is infinite, even for a single hidden node [21]. See also the more recent [5] for complexity bounds on neural networks with piecewise lienar activations. We do not use these kind of activation functions. More useful activation functions are, for example the sigmoid function $\sigma(z)=(1+\exp(z))^{-1}$, whose VC dimensions is bounded by $VCdim(\mathcal{H}_{G(V,E),sigmoid})\leq O\left(|E|^{4}\right)$, or ReLU function, for which $VCdim(\mathcal{H}_{G(V,E),ReLU})\leq O\left(|E|\log|E|l\right)$, with $l$ the network depth. One can limit the capacity by discretizing the weights, e.g. if $w\in[-B,...,B]$, $VCdim\leq 2|E|B$. As we’ve seen, the fact that these network models can have a large capacity is not necessarily good, because capacity comes at the expense of inductive bias.
What about expressivity? Feed-forward neural networks can represent any logical gate, this can be seen as a consequence of [33], and thus any function over $\mathcal{X}=\{\pm 1\}^{d}$ (as proved by Turing [48], see also [20]). Define the class $CIRCUIT_{n}[depth,size]$ as all functions $f:\{\pm 1\}^{n}\rightarrow\{0,1\}$ that can be implemented with at most $size$ AND, OR and NOT gates, and longest path from input to output at most $depth$. We know that circuits can represent any function, see e.g. [2], but only if we are allowed to select an appropriate gate architecture. In neural networks, we keep the architecture fixed (number of nodes, activations, edges) and only vary the weights.
@@
This can be done easily, if we choose the weights of each edge to be ±$\pm 1$ if the edge is connected in the circuit (with /o without a NOT gate in between), 00 otherwise. The bias terms are chosen as $fan\_in-1$ for AND gates, $1-fan\_in$ for OR gates, for a given value of $fan\_in\in[0,1]$. The weights essentially describe which wires exist in the circuit. Thus neural networks can represent any binary function.
More generally, we have a universal representation theorem [17, 25, 3, 41]: Any continuous function $f:[0,1]^{d}\rightarrow\mathbb{R}$ can be approximated to within $\epsilon$ by a feed-forward network with sigmoidal (or almost any other) activation function and a single hidden layer. This shows that as a model class, feed-forward neural nets are extremely expressive and can represent any reality. However, representing functions may require huge networks, e.g. with layer widths exponential in $d$. The relevant question is not what a network of arbitrary size can represent, but what small networks can represent.
@@
Interestingly, a feed-forward neural network can represent any time $T$ computable function with network of size $\tilde{O}(T)$. This is true since anything computable in time $T$ is also computable by a logical circuit of size $\tilde{O}(T)$. This realization has broad implications for machine learning.
Machine learning is an engineering paradigm (of being lazy): use data and examples, instead of expert knowledge and tedious programming, to automatically create efficient systems that solve complex tasks. Therefore, we only care about a model (predictor) $h$ if it can be implemented efficiently. A good learned model only needs to compete with a programmer, producing results that are at least as good as a programmed model in a competitive (in terms of model evaluation) time.
In this case we have a free lunch: the model class $TIME_{T}$ all functions computable by at most time $T$, has capacity $O\left(T\right)$, and hence learnable with $O\left(T\right)$ parameters (e.g. using ERM). Even better: the model class $PROG_{T}$, all functions programmable up to length of code $T$, also has capacity $O\left(T\right)$. This is relatively clear, because the length $T$ bounds the number of bits needed to represent all these functions. Unfortunately, ERM with respect to $PROG_{T}$ is uncomputable. Modified ERM for $TIME_{T}$ (truncating exec. time) can be computed, but is NP-complete. If $P=NP$, we can have universal learning (free lunch). If $P\neq NP$, i.e. if there exist one-way functions that are easy to compute but hard to learn, there is no poly-time algorithm for ERM over $TIME_{T}$. For a more detailed discussion of the representation of functions with circuits, classes of programmable functions and related complexity measures, see e.g. [2].
@@
Going back to neural networks, we completely understand them from a statistical perspective (in terms of capacity and expressivity, see the previous discussion on VC dimension for neural netowrks and universal approximation). The problem with them relates to computation; computing the ERM for feed-forward neural nets is a non-convex optimization problem, and no known algorithm is guaranteed to work. We know that learning in neural nets, even in the simplest cases ($2$-hidden units in one hidden layer), is NP-hard [12]. Even if reality is well-approximated by a small neural net, and we tried optimizing a larger neural net (which has more degrees of freedom), optimization is easy but ERM is still NP-hard. Unfortunately, there is nothing one can do to efficiently solve this computational problem, which is essential to neural nets precisely because of their expressive power : this is the computational no free lunch. Even if a function is exactly representable with single hidden layer and $\Theta(\log d)$ nodes, even with no noise, and even if we take a much larger network or use any other method when learning: no poly-time algorithm can ensure better-than-chance prediction, see e.g. [27, 19].
And nevertheless, deep learning does work! We have seen that from a statistical and computational perspectives, performing ERM on short programs (or short runtime programs) and learning with deep networks is equivalent. Both approaches are universal and approximate reality with reasonable sample complexity. They are both NP-hard, and provably hard to learn with any learning rule (subject to cryptographic assumptions). However there is no practical way to optimize over short programs, as e.g. there is no practical local search over programs. In contrast, deep neural nets are often easy to optimize; they are continuous models, amenable to local search (gradient descent, SGD), and enjoy much empirical success. In the worst case, deep learning is provably impossible, and yet, we are constantly reminded that deep learning is possible. There is a certain property of reality that makes feed-forward neural networks work, especially from the optimization viewpoint, and we have just started to scratch the surface of what that property is. However, we know for sure what it isn’t: it is not the property that reality is well approximated by neural networks.